he proportion of women working in jobs that require analytical skills has been on the rise for several decades. A 2018 study by Pew Research showed that the proportion of women working in jobs that require analytical skills was 0.423. An economist wants to determine if the actual proportion of women working in jobs that require analytical skills is different from the value reported by the Pew Research study.

A normal distribution is used to carry out the following hypothesis test:

H0:p=H0:p= 0.423 vs  HA:p≠HA:p≠ 0.423.

The test statistic for this hypothesis test is 1.7484. Find the p-value. Give your answer to 4 decimal places.
Answer #1

solution

This is the two tailed test .

The null and alternative hypothesis is

H0 : p =0.423

Ha : p 0.423

z=1.7484

P(z > 1.7484) = 1 - P(z < 1.7484) = 1-0.9598=0.0402*2=0.0804

P-value = 0.0804 ( using z table )